OpenStudy (anonymous):

What is the sum of a 12–term arithmetic sequence where the last term is 13 and the common difference is –10?

OpenStudy (anonymous):

There is something wrong @Mimi_x3 ..

OpenStudy (mimi_x3):

I think so too lol I haven't done these in a while..

OpenStudy (mimi_x3):

\[ T_n = a +(n-1)d\] \[ T_{12} = a + (12-1)*-10 = 13\] \[ a-120 +10 = 13\] \[ a -110 = 12\] a = 122

OpenStudy (anonymous):

\[13 = a + (12-1) (-10) \implies 13 = a -110 \implies a = 123\]

OpenStudy (anonymous):

What are we doing we don't know @Mimi_x3 Ha hah a..

OpenStudy (mimi_x3):

Another typo lol \(a = 123\)

OpenStudy (mimi_x3):

Sorry brain dead at 2 am. a -110 = 13 => a =123

OpenStudy (anonymous):

So, \(a = 123\) Now find the sum.. \[\huge Sum = \frac{n}{2}(a + l)\] n = 12. a = 123 and l = 13 Calculate the sum now.. @emkaye16

OpenStudy (anonymous):

What happens?? Another brain dead?? @Mimi_x3

OpenStudy (anonymous):

My internet speed is just as same as that of an ant..

OpenStudy (mimi_x3):

nothing lol i thought there was another mistake; but there wasnt lol

OpenStudy (mimi_x3):

my brain speed is so slow today :(

OpenStudy (anonymous):

@Mimi_x3 I think here we are talking there is no one other watching us.. Ha haha..

OpenStudy (mimi_x3):

lol even the asker is not responding; probably got scared from our mistakes :P

OpenStudy (anonymous):

so the answer is 816?

OpenStudy (anonymous):

Yes you are right @emkaye16 Sorry I was busy somewhere else.. \[\huge \color{blue}{\textbf{Keep It Up...}} \color{green} {\checkmark}\]